OpenStudy (volleyballlover55):

Convert 3,200,000 to scientific notation. 3.2 ⋅ 10^6 3.2 ⋅ 10^5 3.2 ⋅ 10^−6 3.2 ⋅ 10^−5

OpenStudy (volleyballlover55):

@surjithayer

OpenStudy (volleyballlover55):

A?

OpenStudy (anonymous):

correct.

OpenStudy (volleyballlover55):

:)

OpenStudy (volleyballlover55):

Which number is the smallest? 4.4 ⋅ 10^−5, 3.99 ⋅ 10^−6, 2.6 ⋅ 10^2, 1.85 ⋅ 10^3 4.4 ⋅ 10^−5 3.99 ⋅ 10^−6 2.6 ⋅ 10^2 1.85 ⋅ 10^3

OpenStudy (anonymous):

\[4.4*10^{-5}=\frac{ 44 }{ 10 }*10^{-5}=44*10^{-6}\] and 3.9*10^{-6} which is smaller.

OpenStudy (anonymous):

out of 44 and 3.9 which is smaller.

OpenStudy (volleyballlover55):

3.9

OpenStudy (anonymous):

obviously it is smallest

OpenStudy (volleyballlover55):

ok so then the answer is b?

OpenStudy (anonymous):

correct.

OpenStudy (volleyballlover55):

;)

OpenStudy (volleyballlover55):

Convert 0.0001 to a power of 10. 10−4 104 10−3 103

OpenStudy (volleyballlover55):

A?